The boy pulls a sledge weighing 5 kg on the snow with the help of a rope that builds an angle of 60 ° with the horizontal direction. The force with which the boy pulls the rope is 10 N, and the sled moves at a constant speed. Determine the coefficient of friction between the sled and the snow surface.


1
Expert's answer
2021-03-24T19:43:37-0400

The force of friction is


"f=\mu N."


The normal force N can be found from the following condition:


"N-mg+F\sin60°=0,\\\space\\
\mu=\frac fN=\frac{f}{mg-F\sin60°}."

Since the sled moves at a constant velocity, we have the following equilibrium of the pulling force and the friction:


"f-F\cos60°=0,\\
f=F\cos60°."

Therefore:


"\mu=\frac{F\cos60°}{mg-F\sin60°}=0.124."